OpenStudy (anonymous):

Multiply and simplify the product: (2sqrt7 - sqrt3)^2

OpenStudy (anonymous):

it's simply (a-b)^2 = a^2 -2*a*b + b^2.

OpenStudy (anonymous):

Is the anwer 31-2sqrt21??

OpenStudy (anonymous):

i think it would be 4sqrt21 instead of 2

OpenStudy (anonymous):

yeah it should be \[31-4\sqrt{21}\]

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

the middle term is 2*a*b, and a*b = 2rt7*rt3 = 2rt21 so 2*a*b = 2*2rt21